# Generalizations of Koga’s version of the Wiener-Ikehara theorem

### v2

@@ 1. Introduction
The Wiener-Ikehara theorem [37] is a foundational result in complex Tauberian theory. Originally devised to significantly simplify an early result of Landau [26] and so deliver one of the quickest deductions of the prime number theorem, it has found countless applications in diverse areas of mathematics such as operator theory, partial differential equations, and number theory. The interested reader is referred to the books [13, Chapter 10], [24, Chapter III], and [34, Chapter II.7] for excellent accounts on the subject and the recent articles [5, 10, 12, 20, 22, 30, 38] for some developments during the last decade. See also [1, 6, 9, 11, 33] for closely related complex Tauberian theorems for Laplace transforms, such as the Ingham-Karamata theorem.
@@
Naturally, the hypothesis of $L^{1}_{loc}$-boundary behavior covers the case of continuous extension, and in particular that of analytic continuation. On the other hand, we point out that the boundary requirements on the Laplace transform can further be taken to a minimum if one employs the so-called local pseudofunction boundary behavior (cf. [10, 12, 25]). The pseudofunction approach plays a major role in modern complex Tauberian theory (cf. [24, Chapter III]).
@@
Koga’s main motivation to establish a novel version of the Wiener-Ikehara theorem was to provide a Dirichlet series generalization of the Kolmogorov-Erdős-Feller-Pollard renewal theorem [14, 23] (cf. [18, Sections XIII.3 and XIII.11]). Moreover, he also obtained a Tauberian theorem for power series and applied it to give a new proof of the classical quoted renewal theorem. Upon a minor reformulation (cf. Remark 7.2), Koga’s Tauberian theorem for Laplace transforms reads:
Theorem 1.1 ([22, Theorem 2]).
@@
The aim of this paper is to extend Koga’s theorem by further considering other useful hypotheses for the boundary behavior of $\Re e\>\mathcal{L}\{S;s\}$ near $s=1$. In addition, our results considerably improve Koga’s ones, in the sense that we have been able to replace the somehow unnatural condition (1.3) (and its analog for power series) by just requirements on the Laplace transform. Notice that (1.3) always implies (1.7) below with $k=2$, so that the next theorem, our first main result, includes Theorem 1.1 as its particular instance, and, more generally, also shows that one can relax (1.3) in Theorem 1.1 to
(1.6) $\int_{1}^{\infty}\frac{|S(x)|}{x^{k}e^{x}}\mathrm{d}x<\infty\qquad\mbox{for some }k\in\mathbb{N};$
see also Remark 7.1. (It is worth noting that these integral growth conditions could still be significantly weakened, see (7.6) and (7.7) in Remark 7.4.)
@@
Let $S\in L^{1}_{loc}[0,\infty)$ be log-linearly slowly decreasing and have convergent Laplace transform on $\Re e\>s>1.$ Suppose that the harmonic function $U(s)=\Re e\>\mathcal{L}\{S;s\}$ has $L^{1}_{loc}$-boundary behavior on $1+i(\mathbb{R}\setminus\{0\})$ and that there is some $\lambda>0$ such that one of the following three conditions holds:
@@
    there is $c\geq 0$ such that $U(\sigma+it)\geq-c$, for $t\in(-\lambda,\lambda)$ and $\sigma\in(1,2]$;
@@
    $\sup_{1<\sigma<2}\int_{-\lambda}^{\lambda}|U(\sigma+it)|\mathrm{d}t<\infty$;
  • (B.3)
    there are $g\in L^{1}(-\lambda,\lambda)$ and $k\in\mathbb{N}$ such that (1.4) holds and
    (1.7) $U(\sigma+it)=O\left({(\sigma-1)^{-k}}\right),\qquad\mbox{for }t\in(-\lambda,\lambda)\mbox{ and }\sigma\in(1,2].$
@@
Working with this new formulation has great practical value as in certain situations it is easier to apply than Theorem 1.2. In fact, we will deduce the following corollary of Ingham-Karamata type from Theorem 1.3. We shall exemplify its usefulness in Section 2 by giving a quick Tauberian proof of Blackwell’s renewal theorem [3]. We also give there a simpler treatment of Koga’s renewal theorem for Dirichlet series [22, Theorem 5] based on Theorem 1.3.
@@
Let $T$ be a function on $[0,\infty)$ such that $T(x)+Mx$ is non-decreasing, for some constant $M>0$, and such that its Laplace-Stieltjes transform converges on $\Re e\>s>0$. Suppose that $U(s)=\Re e\>\int_{0^{-}}^{\infty}e^{-sx}\mathrm{d}T(x)$ has $L^{1}_{loc}$-boundary behavior on $i\mathbb{R}\setminus\{0\}$ and that there is some $\lambda>0$ such that one of the following three conditions holds:
@@
    there is $c\geq 0$ such that $U(\sigma+it)\geq-c$, for $t\in(-\lambda,\lambda)$ and $\sigma\in(0,1]$;
@@
  • (B0.3)
    there are $g\in L^{1}(-\lambda,\lambda)$ and $k\in\mathbb{N}$ such that $U(\sigma+it)\geq g(t),$ for a.e. $t\in(-\lambda,\lambda)$ and $\sigma\in(0,1]$, and
    $U(\sigma+it)=O\left({\sigma^{-k}}\right),\qquad\mbox{for }t\in(-\lambda,\lambda)\mbox{ and }\sigma\in(0,1].$
@@
We shall also prove the next Tauberian theorem for power series, which improves upon [22, Theorem 3].
@@
Let $F(z)=\sum_{n=0}^{\infty}c_{n}z^{n}$ be analytic on the unit disc $\mathbb{D}$ with real coefficients $\{c_{n}\}_{n=0}^{\infty}$. Suppose that the harmonic function $U(z)=\Re e\>F(z)$ has $L^{1}_{loc}$-boundary behavior on $\partial\mathbb{D}\setminus\{1\}$ and there is some $\theta_{0}\in(0,\pi)$ such that one of the following three conditions holds:
@@
    there is $c\geq 0$ such that $U(re^{i\theta})\geq-c,$ for $\theta\in(-\theta_{0},\theta_{0})$ and $r\in[0,1)$;
@@
  • (b.3)
    there are $g\in L^{1}(-\theta_{0},\theta_{0})$ and $k\in\mathbb{N}$ such that
    (1.12) $U(re^{i\theta})=O\left({(1-r)^{-k}}\right),\qquad\mbox{for }\theta\in(-\theta_{0},\theta_{0})\mbox{ and }r\in[0,1)$
    and
    (1.13) $U(re^{i\theta})\geq g(\theta),\qquad\mbox{for a.e. }\theta\in(-\theta_{0},\theta_{0})\mbox{ and }r\in[0,1).$
@@
The plan of the article is as follows. We discuss in Section 2 how Theorem 1.3 and Theorem 1.5 can be applied to renewal theory [19, Chapter XI]; our applications emphasize the role of the assumptions (B.1), (B0.1), and (b.1) in the corresponding cases, which make the theorems relatively simple to apply. In Section 3 we obtain a slight extension of the exact Wiener-Ikehara Tauberian theorem [10, Theorem 3.6], where we shall show that (1.2) holds if and only if $S$ is log-linearly slowly decreasing, its Laplace transform converges for $\Re e\>s>1$, and the real part of the function $G$ given by (1.1) has so-called local pseudofunction boundary behavior on $1+i\mathbb{R}$. Section 4 is devoted to the proofs of Theorem 1.2 and Theorem 1.3; our approach there will be to reduce them to the exact Wiener-Ikehara theorem from Section 3. Theorem 1.5 will be shown in Section 5, while a proof of Corollary 1.4 will be given in Section 6. Finally, we close the article with some remarks and further extensions of our Tauberian theorems, which will be discussed in Section 7.
@@ 2. Application: Renewal theorems
Since we assumed $\alpha$ to be maximal, we have that $1=\mathrm{gcd}\{n:p_{n}\neq 0\}$, which implies that $G(re^{i\theta})\neq 1$ for all $\theta\in[-\pi,\pi]\setminus\{0\}$. Here $G$ stands for the power series $G(z)=\sum_{n=1}^{\infty}p_{n}z^{n}$, which is continuous on the closed unit disc. Due to (2.4), we obtain $F(z)=\sum_{n=0}^{\infty}q_{n}z^{n}=(1-G(z))^{-1}$. As in the previous the case, we also have $\Re e\>F(z)>0$ for all $z\in\mathbb{D}$ (since $\Re e\>G(z)<1$ on $\mathbb{D}$). Hence, Theorem 1.5 allows us to conclude that
@@
We set $S(x)=\sum_{n\leq e^{x}}f(n)$, so that $F(s)=\mathcal{L}\{\mathrm{d}S;s\}=\sum_{n=1}^{\infty}f(n)/n^{s}$. The familiar properties of Dirichlet series and $f=e+f\star g$ yield (2.3) with now $G$ given by the Dirichlet series of $g$, i.e., $G(s)=\sum_{n=2}^{\infty}g(n)/n^{s}$, which continuously extends to $\Re e\>s=1$. In view of [22, Lemma 11], the assumption on $\{n:\>g(n)\neq 0\}$ implies that $G(1+it)\neq 1$ for $t\neq 0$. Obviously, $\Re e\>F(s)>0$ on the entire open half-plane $\Re e\>s>1$. An application of Theorem 1.3 thus shows that
@@ 3. Exact Wiener-Ikehara theorem revisited
$U(\sigma+it)=O((\sigma-1)^{-k})$
@@
By [10, Proposition 2.1], $\langle f(t),\varphi(t)e^{iht}\rangle=o(1)$ as $h\to-\infty$, for each $\varphi\in\mathcal{D}(I)$ and our assumption is $\langle u(t),\varphi(t)e^{iht}\rangle=o(1)$ as $|h|\to\infty$. We also notice that $u$ is an even distribution, while $f(t)-u(t)$ is odd. For $h>0$ and $\varphi\in\mathcal{D}(I)$ real-valued and even,
@@
Lemma 3.1 and [10, Theorem 3.6] together thus yield:
@@
Lemma 3.1 highlights a key advantage of the local pseudofunction approach over $L^{1}_{loc}$-boundary behavior. In fact, it is well-known that if a harmonic function has $L^{1}_{loc}$-boundary behavior, the distributional boundary values of its harmonic conjugate functions do not necessarily belong to $L^{1}_{loc}$ (see e.g. [21, p. 73])
@@
We first claim that any of our assumptions imply that $U$ admits a boundary distribution on $1+i\mathbb{R}$. This is actually our hypothesis away from the boundary point $s=1$, where we even have the stronger $L_{loc}^{1}$-boundary behavior. So, we must then still establish the existence of a boundary distribution in a boundary neighborhood of 1. We notice that the bound (1.7) is actually equivalent to it (cf. Section 3). Let us verify that (1.7) also remains valid under either assumption (B.1) or (B.2). Let $\Phi:\mathbb{D}\to\Omega$ be a conformal equivalence between the unit disc and a region $\Omega\subset\{s:\>\Re e\>s>1\}$ whose boundary is a smooth Jordan curve that meets the line $\Re e\>s=1$ in a closed segment containing $s=1$ and inside the interval $1+i(-\lambda,\lambda)$. Set $V=U\circ\Phi$. If (B.1) holds, then $V(z)+c$ is a non-negative harmonic function. On the other hand, if (B.2) is satisfied, $V$ belongs to the harmonic Hardy space $h^{1}(\mathbb{D})$, as inferred from [^4] Theorem 10.1 is only stated for analytic functions in [15, p. 168], but the proof given there applies to harmonic functions as well. [15, Theorem 10.1, p. 168] and [29, Theorem 3.5, p. 48]. Hence [15, Theorem 1.1, p. 2], $V$ is a Poisson-Stieltjes integral, whence we readily obtain the bound
@@
Since $\Phi$ extends to a diffeomorphism [^5] Actually, conformally extends to a complex neighborhood of $1$ as an application of the Schwarz reflection principle for analytic arcs yields. (by [29, Theorem 3.5, p. 48] again) between complex neighborhoods of an arc of $\partial\mathbb{D}$ and a segment $1+iI$ containing $1$, we also have
$U(\sigma+it)=O\left(\frac{1}{\sigma-1}\right),\qquad\sigma+it\in(1,2]\times I.$
Let now $u(t)=\lim_{\sigma\to 1^{+}}U(\sigma+it)$ in $\mathcal{D}^{\prime}(\mathbb{R})$. By assumption $u=f$ on $\mathbb{R}\setminus\{0\}$ with $f\in L^{1}_{loc}(\mathbb{R}\setminus\{0\})$. If either (B.1) or (B.3) hold, then $u-g$ should be a non-negative measure on $(-\lambda,\lambda)$, where we set $g=-c$ in the case of condition (B.1). When (B.2) holds, $u$ is a measure on $(-\lambda,\lambda)$, as follows from the Banach-Alaoglu theorem (or Helly’s selection principle as better known in the Lebesgue-Stieltjes measure context), because (B.2) tells us that $\{U(\sigma+i\>\cdot\>)\>:1<\sigma<2\}$ is bounded in the dual of $C[-\lambda,\lambda]$. Summarizing, in every case we have shown that the distribution $u$ is a Radon measure on $\mathbb{R}$. Using the Lebesgue decomposition of $u$ (Lebesgue-Radon-Nikodym theorem [32, p. 121]), we conclude that $f\in L^{1}_{loc}(\mathbb{R})$ and that $f$ is its absolutely continuos part, while its singular part must have point support at 0. Hence, $u=f+\pi a\delta$ for some constant $a\in\mathbb{R}$, where as usual $\delta$ stands for the Dirac delta distribution. Finally, using the well-known formula (cf. [17, Eq. (2.17), p. 58])
@@
$h\to 0^{+}$. The hypotheses (B0.1)–(B0.3) translate into the conditions (B.1)–(B.3) for the real part of the Laplace-Stieljes transform of $S$. Theorem 1.3 then yields $S(x)\sim ae^{x}$ with $a$ given by (1.11). Writing $\tau(x)=e^{-x}S(x)-a=o(1)$, noticing that $\mathrm{d}Q(x)=e^{-x}\mathrm{d}S(x)$, and integrating by parts, we obtain,
@@
7. Further extensions and concluding remarks

### v3

@@ 1. Introduction
The Wiener-Ikehara theorem [32] is a foundational result in complex Tauberian theory. Originally devised to significantly simplify an early result of Landau [22] and so deliver one of the quickest deductions of the prime number theorem, it has found countless applications in diverse areas of mathematics such as operator theory, partial differential equations, and number theory. The interested reader is referred to the books [8, Chapter 10], [19, Chapter III], and [30, Chapter II.7] for excellent accounts on the subject and the recent articles [3, 6, 7, 15, 17, 24, 26, 33] for some developments during the last decade. See also [1, 4, 21, 29] for closely related complex Tauberian theorems for Laplace transforms, such as the Ingham-Karamata theorem.
@@
Naturally, the hypothesis of $L^{1}_{loc}$-boundary behavior covers the case of continuous extension, and in particular that of analytic continuation. On the other hand, we point out that the boundary requirements on the Laplace transform can further be taken to a minimum if one employs the so-called local pseudofunction boundary behavior (cf. [6, 7, 20]). The pseudofunction approach plays a major role in modern complex Tauberian theory (cf. [19, Chapter III]).
@@
Koga’s main motivation to establish a novel version of the Wiener-Ikehara theorem was to provide a Dirichlet series generalization of the Kolmogorov-Erdős-Feller-Pollard renewal theorem [9, 18] (cf. [13, Sections XIII.3 and XIII.11]). Moreover, he also obtained a Tauberian theorem for power series and applied it to give a new proof of the classical quoted renewal theorem. Upon a minor reformulation (cf. Remark 7.2), Koga’s Tauberian theorem for Laplace transforms reads:
Theorem 1.1 ([17, Theorem 2]).
@@
The aim of this paper is to considerably improve Koga’s theorem by showing that it still holds true if one removes condition (1.3) from its set of hypotheses. In addition to hold under weaker assumptions, we shall also consider a new useful alternative hypothesiss for the boundary behavior of $\Re e\>\mathcal{L}\{S;s\}$ near $s=1$.
@@
Let $S\in L^{1}_{loc}[0,\infty)$ be log-linearly slowly decreasing and have convergent Laplace transform on $\Re e\>s>1.$ Suppose that the harmonic function $U(s)=\Re e\>\mathcal{L}\{S;s\}$ has $L^{1}_{loc}$-boundary behavior on $1+i(\mathbb{R}\setminus\{0\})$ and that there is some $\lambda>0$ such that one of the following two conditions holds:
@@
    there is $g\in L^{1}(-\lambda,\lambda)$ such that (1.4) holds;
@@
    $\sup_{1<\sigma<2}\int_{-\lambda}^{\lambda}|U(\sigma+it)|\mathrm{d}t<\infty$.
@@
Working with this new formulation has great practical value as in certain situations it is easier to apply than Theorem 1.2. In fact, we will deduce the following corollary of Ingham-Karamata type from Theorem 1.3. We shall exemplify its usefulness in Section 2 by giving a quick Tauberian proof of Blackwell’s renewal theorem [2]. We also give there a simpler treatment of Koga’s renewal theorem for Dirichlet series [17, Theorem 5] based on Theorem 1.3.
@@
Let $T$ be a function on $[0,\infty)$ such that $T(x)+Mx$ is non-decreasing, for some constant $M>0$, and such that its Laplace-Stieltjes transform converges on $\Re e\>s>0$. Suppose that $U(s)=\Re e\>\int_{0^{-}}^{\infty}e^{-sx}\mathrm{d}T(x)$ has $L^{1}_{loc}$-boundary behavior on $i\mathbb{R}\setminus\{0\}$ and that there is some $\lambda>0$ such that one of the following two conditions holds:
@@
    there is $g\in L^{1}(-\lambda,\lambda)$ and $k\in\mathbb{N}$ such that $U(\sigma+it)\geq g(t)$ for a.e. $t\in(-\lambda,\lambda)$ and $\sigma\in(0,1]$;
@@
We shall also prove the next Tauberian theorem for power series, which improves upon [17, Theorem 3].
@@
Let $F(z)=\sum_{n=0}^{\infty}c_{n}z^{n}$ be analytic on the unit disc $\mathbb{D}$ with real coefficients $\{c_{n}\}_{n=0}^{\infty}$. Suppose that the harmonic function $U(z)=\Re e\>F(z)$ has $L^{1}_{loc}$-boundary behavior on $\partial\mathbb{D}\setminus\{1\}$ and there is some $\theta_{0}\in(0,\pi)$ such that one of the following two conditions holds:
@@
    there is $g\in L^{1}(-\theta_{0},\theta_{0})$ such that $U(re^{i\theta})\geq g(\theta)$ for a.e. $\theta\in(-\theta_{0},\theta_{0})$ and $r\in[0,1)$;
@@
The plan of the article is as follows. We discuss in Section 2 how Theorem 1.3, Corollary 1.4, and Theorem 1.5 can be applied to renewal theory [14, Chapter XI]; our applications emphasize the role of the assumptions (B.1), (B0.1), and (b.1) in the corresponding cases, which make the theorems relatively simple to apply. In Section 3, we obtain a slight extension of the exact Wiener-Ikehara Tauberian theorem [6, Theorem 3.6], where we shall show that (1.2) holds if and only if $S$ is log-linearly slowly decreasing, its Laplace transform converges for $\Re e\>s>1$, and the real part of the function $G$ given by (1.1) has so-called local pseudofunction boundary behavior on $1+i\mathbb{R}$. Section 4 is devoted to the proofs of Theorem 1.2 and Theorem 1.3; our approach there will be to reduce them to the exact Wiener-Ikehara theorem from Section 3. Theorem 1.5 will be shown in Section 5, while a proof of Corollary 1.4 will be given in Section 6. Finally, we close the article with some remarks and further extensions of our Tauberian theorems, which will be discussed in Section 7.
@@ 2. Application: Renewal theorems
Since we assumed $\alpha$ to be maximal, we have $1=\mathrm{gcd}\{n:p_{n}\neq 0\}$, which implies that $G(re^{i\theta})\neq 1$ for all $\theta\in[-\pi,\pi]\setminus\{0\}$. Here $G$ stands for the power series $G(z)=\sum_{n=1}^{\infty}p_{n}z^{n}$, which is continuous on the closed unit disc. Due to (2.4), we obtain $F(z)=\sum_{n=0}^{\infty}q_{n}z^{n}=(1-G(z))^{-1}$. As in the previous the case, we also have $\Re e\>F(z)>0$ for all $z\in\mathbb{D}$ (since $\Re e\>G(z)<1$ on $\mathbb{D}$). Hence, Theorem 1.5 allows us to conclude that
@@
We set $S(x)=\sum_{n\leq e^{x}}f(n)$, so that $F(s)=\mathcal{L}\{\mathrm{d}S;s\}=\sum_{n=1}^{\infty}f(n)/n^{s}$. The familiar properties of Dirichlet series and $f=e+f\star g$ yield (2.3) with now $G$ given by the Dirichlet series of $g$, i.e., $G(s)=\sum_{n=2}^{\infty}g(n)/n^{s}$, which continuously extends to $\Re e\>s=1$. In view of [17, Lemma 11], the assumption on $\{n:\>g(n)\neq 0\}$ implies that $G(1+it)\neq 1$ for $t\neq 0$. Obviously, $\Re e\>F(s)>0$ on the entire open half-plane $\Re e\>s>1$. An application of Theorem 1.3 thus shows that
@@ 3. Exact Wiener-Ikehara theorem revisited
(3.1) $U(\sigma+it)=O\left(\frac{1}{(\sigma-1)^{k}}\right)$
@@
By [6, Proposition 2.1], $\langle f(t),\varphi(t)e^{iht}\rangle=o(1)$ as $h\to-\infty$, for each $\varphi\in\mathcal{D}(I)$ and our assumption is $\langle u(t),\varphi(t)e^{iht}\rangle=o(1)$ as $|h|\to\infty$. We also notice that $u$ is an even distribution, while $f(t)-u(t)$ is odd. For $h>0$ and $\varphi\in\mathcal{D}(I)$ real-valued and even,
@@
Lemma 3.1 and [6, Theorem 3.6] together thus yield:
@@
Lemma 3.1 highlights a key advantage of the local pseudofunction approach over $L^{1}_{loc}$-boundary behavior. In fact, it is well-known that if a harmonic function has $L^{1}_{loc}$-boundary behavior, the distributional boundary values of its harmonic conjugate functions do not necessarily belong to $L^{1}_{loc}$ (see e.g. [16, p. 73])
@@
The first part of our proof consists in showing that any of our assumptions imply that $U$ admits a boundary distribution on $1+i\mathbb{R}$. This is actually our hypothesis away from the boundary point $s=1$, where we even have the stronger $L_{loc}^{1}$-boundary behavior. So, we must then still establish the existence of a boundary distribution in a boundary neighborhood of 1. We need some preparation for it.
Let $\Phi:\mathbb{D}\to\Omega$ be a conformal equivalence between the unit disc and a region $\Omega\subset\{s:\>\Re e\>s>1\}$ whose boundary is a smooth (namely, $C^{\infty}$) Jordan curve that meets the line $\Re e\>s=1$ in a closed segment containing the interval $1+i(-\lambda,\lambda)$. Classical results from the theory of conformal maps (see [25, Theorem 2.6, p. 24; Theorem 3.5, p. 48; Theorem 3.6, p. 49] guarantee that $\Phi$ extends to a smooth diffeomorphism $\Phi:\overline{\mathbb{D}}\to\overline{\Omega}$, and we may assume that $\Phi(1)=1.$ Moreover, if $\Phi(J)=1+i[-\lambda,\lambda]$, then $\Phi$ has analytic extension through an open circular arc containing $J$, as one infers from the Schwarz reflection principle for analytic arcs. We consider the harmonic function $V=U\circ\Phi$.
Claim 4.1.
The function $V$ belongs to the harmonic Hardy space $h^{1}(\mathbb{D})$.
Proof.
If (B.2) is satisfied, we directly get $V\in h^{1}(\mathbb{D})$, as inferred from [^4] Theorem 10.1 is only stated for analytic functions in [10, p. 168], but the proof given there applies to harmonic functions as well. [10, Theorem 10.1, p. 168] and [25, Theorem 3.5, p. 48]. Assume now that (B.1) holds. Applying again [10, Theorem 10.1, p. 168], we conclude that $V$ has $L^{1}_{loc}$-boundary behavior on any boundary arc that does not contain the boundary point $1$. Combining the latter fact with a.e. existence of the radial boundary limits of $V$, we obtain that $V$ belongs to the harmonic Hardy space $h^{1}(S)$ for any disc sector of the form
$S=\{z\in\mathbb{D}:\>\arg z\notin J^{\prime}\}\qquad\mbox{with a closed subarc}1\in J^{\prime}\subsetneq J.$
Let us fix such a sector $S$. We also write $J=\{e^{i\theta}:\theta\in[\alpha,\beta]\}$. We have
$\sup_{0<r<1}\int_{[-\pi,\pi]\setminus{(\alpha,\beta)}}|V(re^{i\theta})|\mathrm{d}\theta<\infty,$
because $V\in h^{1}(S)$. Therefore, we just need to bound the integral of $|V(re^{i\theta})|$ over $[\alpha,\beta]$ as $r\to 1^{-}$. Using the mean value property of harmonic functions and condition (B.1), we obtain
$\int_{\alpha}^{\beta}|V(re^{i\theta})|\mathrm{d\theta}$ $\leq\int_{\alpha}^{\beta}|g(\Im m\>(\Phi(re^{i\theta})))|\mathrm{d}\theta+\int_{\alpha}^{\beta}[V(re^{i\theta})-g(\Im m\>(\Phi(re^{i\theta})))]\mathrm{d}\theta$
$\leq 2\int_{\alpha}^{\beta}|g(\Im m\>(\Phi(re^{i\theta})))|\mathrm{d}\theta+2\pi V(0)+\int_{[-\pi,\pi]\setminus{(\alpha,\beta)}}|V(re^{i\theta})|\mathrm{d}\theta$
$=2\int_{\alpha}^{\beta}|g(\Im m\>(\Phi(e^{i\theta})))|\mathrm{d}\theta+o(1)+2\pi V(0)+\int_{[-\pi,\pi]\setminus{(\alpha,\beta)}}|V(re^{i\theta})|\mathrm{d}\theta$
$=O(1),\qquad r\to 1^{-},$
where in the third line we have used that $\Phi(re^{i\theta})\to\Phi(ie^{\theta})$ and $(\Phi(re^{i\theta}))^{\prime}\to(\Phi(ie^{\theta}))^{\prime}$ uniformly on $[\alpha,\beta]$. This shows that $V\in h^{1}(\mathbb{D})$, also under (B.1). ∎
We can now establish our original claim:
Claim 4.2.
$U$ admits a boundary distribution on a boundary neighborhood of 1.
Proof.
In view of [10, Theorem 1.1, p. 2] and Claim 4.1, the harmonic function $V$ is a Poisson-Stieltjes integral, whence we readily obtain the bound
@@
Since $\Phi$ extends to a diffeomorphism between complex neighborhoods of $J$ and $1+i[-\lambda,\lambda]$, we also have
$U(\sigma+it)=O\left(\frac{1}{\sigma-1}\right),\qquad\sigma+it\in(1,2]\times[-\lambda,\lambda].$
The latter bound yields the claim (see Section 3). ∎
We can now move to the second part of proof. Let
$u(t)=\lim_{\sigma\to 1^{+}}U(\sigma+it)\qquad\mbox{in }\mathcal{D}^{\prime}(\mathbb{R}).$
By assumption $u=f$ on $\mathbb{R}\setminus\{0\}$ with $f\in L^{1}_{loc}(\mathbb{R}\setminus\{0\})$. If (B.1) holds, then $u-g$ should be a non-negative measure on $(-\lambda,\lambda)$. When (B.2) is satisfied, $u$ is a measure on $(-\lambda,\lambda)$, as follows from the Banach-Alaoglu theorem (or Helly’s selection principle as better known in the Lebesgue-Stieltjes measure context), because (B.2) tells us that $\{U(\sigma+i\>\cdot\>)\>:1<\sigma<2\}$ is bounded in the dual of $C[-\lambda,\lambda]$. Summarizing, in every case we have shown that the distribution $u$ is a Radon measure on $\mathbb{R}$. Using the Lebesgue decomposition of $u$ (Lebesgue-Radon-Nikodym theorem [28, p. 121]), we conclude that $f\in L^{1}_{loc}(\mathbb{R})$ and that $f$ is its absolutely continuos part, while its singular part must have point support at 0. Hence, $u=f+\pi a\delta$ for some constant $a\in\mathbb{R}$, where as usual $\delta$ stands for the Dirac delta distribution. Finally, using the well-known formula (cf. [12, Eq. (2.17), p. 58])
@@
$h\to 0^{+}$. The hypotheses (B0.1) and (B0.2) translate into the conditions (B.1) and (B.2), respectively, for the real part of the Laplace-Stieljes transform of $S$. Theorem 1.3 then yields $S(x)\sim ae^{x}$ with $a$ given by (1.9). Writing $\tau(x)=e^{-x}S(x)-a=o(1)$, noticing that $\mathrm{d}Q(x)=e^{-x}\mathrm{d}S(x)$, and integrating by parts, we obtain,
@@
7. Concluding remarks